Author Archives: shetrone

The week of November 6

A brief advertisement before we get into this week’s news: One of our Resident Astronomers is retiring. If you are a PhD astronomer and interested in working on one of the largest optical telescopes in the world doing cutting edge research please consider applying: https://utdirect.utexas.edu/apps/hr/jobs/nlogon/171020014209

This week the big changes to the telescope have been the removal of the Prime Focus Instrument Platform (PFIP) work platform and the arrival of 3 more VIRUS units. The PFIP work platform, as the name suggests, allowed us to drop people off to work on the PFIP. We will still be able to work on the PFIP but now we have to do it from the JLG worker-lift. The work platform covered up a number of mirrors (about 10% of the light collecting area) as can be seen in the images below taken with our pupil viewing camera.

Two images of the HET pupil taken with our pupil viewing camera before (left) and after (right) the PFIP work platform was removed.

With the delivery of 3 more VIRUS units we now have 28 working VIRUS units or 56 spectrographs or more than 12,000 fibers on the sky!

The week of October 16

The exciting news this week has been the arrival of the Habitable Planet Finder (HPF). This is the first of our new high resolution instruments and an instrument well suited to working in bright moon conditions. The HPF was designed and built by our Penn State partners and arrived on the 16th.

This instrument is designed for extremely high precision spectroscopy capable of detecting the reflex motion of stars as small earth sized planets go around them. To achieve that precision the instrument is housed in our temperature controlled basement at the HET inside a temperature controlled room inside a large temperature controlled vacuum chamber. All of these efforts allow them to control the temperature of the optics of the instrument at a level of 0.001 degrees Celsius.

After very carefully cleaning the enclosure that will house the vacuum chambered instrument the HPF team

was able to open their instrument and after a very through inspection proudly announced that they have just as many pieces of glass as they did in the assembly lab at Penn State (an optics joke). After a few final checks and the inclusion of their single moving part inside of the spectrograph they sealed up the vacuum chamber which, if things continue to go very well, may remain sealed for several years to come. The process of pumping the vacuum out of the large chamber took the rest of the weekend.

In the coming days and weeks the HPF team will monitor its stability, install the laser metrology system and get the systems ready for on-sky commissioning.

Quite an exhausting and exciting week at the HET!

The week of October 2nd

This week we are pleased to announce that a new VIRUS unit was installed in side two of the VIRUS enclosure. This brings us to 22 VIRUS units or 44 spectrographs. We also took a little time in the last engineering run to add on some valves to the vacuum fittings which will allow us to cold pump on the VIRUS units which takes far less time to do than to warm up and then repump which was our older methodology. Keeping 22 VIRUS units going is starting to be a little easier but still takes a lot of management.

In addition to the work on VIRUS, we have also installed in the coherent fiber bundles for the HPF. These coherent fiber bundles will be used to setup stars on HPF science fibers. HPF will arrive in the coming weeks and we are very excited to get our first high resolution instrument on sky in the coming months.

The week of September 18

We are in another 3 week science period but our big news this week is that we have been able to bring a few more VIRUS Units on-line which brings our total number of active units up to 21 units. Recall that a VIRUS unit is made up of 2 spectrographs so we now have 42 spectrographs on-line. These newest units are located in VIRUS enclosure 2 which means that we are now making use of both of the “saddle bags” which were installed at the HET as part of the Wide Field Upgrade. This makes the VIRUS closer to being ready for the main part of the HETDEX survey. As we add more Spectrograph units the system continues to get more complex with multiplexers and timing systems.

In some recent very clear nights we were able to observe 10 HETDEX shots (what we call a pointing for the survey) in a single night. This is a new record and our setup times are now typically below 5 minutes when moving from one shot to another. Further small milestones….

The week of September 4th

We have now observed one month out of four in the 2017-3 period. The month started off pretty slow with lots of bad weather in August but we have had some good clear nights. We are still conducting 7 days of engineering around full moon, at least until we get our first bright time instrument. So far we have collected 52.9 hours of charged time and completed 21.8% of the TAC allocated time. We have started collecting the first LRS2 Guaranteed Time Observations (GTO) allocated by the HET board for the instrument team as well as the first time allocated by the HET board for the HETDEX experiment. The latter includes time on well studied fields to make sure that our Lyman alpha emitters detected not really other types of emission sources and to better quantify our sensitivity/throughput.

The week of August 7

We have completed the 2017-2 trimester (April – July). It was a very successful. The HET Board allowed us to transition from 2 weeks of science operations per lunation to 3 weeks of science operations. We also completed a large commissioning effort of 27 hours of VIRUS exposures on well studied EGS fields. Below are some statistics from the trimester: 138.5 hours of acceptable high priority exposures and 40.5 hours of low priority filler exposures. The average overhead for LRS2 per requested visit was 7.9 minutes which is nearly a factor of two better than the overheads with LRS before the Wide Field Upgrade. The average visit length was 18 minutes long and the longest completed visit was 120 minutes. The completion rate for priority 0 targets was 97.2%, for priority 1 targets was 98.1%, for priority 2 targets was 90.1% and for priority 3 targets was 96.4%.

For the new 2017-3 trimester (August – November) we will continue with the 3 weeks of science per lunation. The HET board has approved the first HETDEX time as well as Guaranteed Time Observing (GTO) for the commissioning team of LRS2. In addition, we are expecting to see the arrival of the Habitable-zone Planet Finder (HPF) spectrograph.

The week of February 27

This week marked the end of another dark run and thus the end of another science period. We are now also three quarters of the way through the current science period labeled 17-1. As such we are starting to see a larger number of filler Priority 4 targets being observed. In this last dark run we observed 374 acceptable observations for a total of 38 hours and 122 Priority 4 targets for a total of 22 hours. We saw 6% of the time lost due to problems. This is a fairly respectable number and fairly typical of the mature HET before the takedown.

We are now three quarters of the way through the current 17-1 period. The Telescope Allocation Committees have begun meeting and the new allocations are being submitted for the 17-2 period which begins April 1. For this period the same suite of instruments will be available for the PIs: LRS2-B and LRS2-R and the two weeks around dark time will also be fully committed to science observing.

The week of February 13

Despite some pretty poor weather this past week we are finished with the bright time engineering run and ready to start science over the 14 days around new moon. In this past engineering run we found some problems with the wave-front sensor software and implemented a new forward model of the hexapod motion. This new model allows for us to compensate for an unintended very small rotation that is induced by hexapod motion over the course of a trajectory. The impact of this rotation was to cause the field to rotate by several arcseconds over a trajectory. Our new forward model eliminates all of the hexapod induced rotation but allowed us to see a much smaller Azimuth dependent rotation term. This much smaller term will be addressed after this next science run, during the next full moon. Overall, a good improvement to the telescope’s operation.

The Week of January 30th

The end of this week marks us as half way through the current 4 month period (December – March). We have completed 70% of the highest priority band P1, 52% of the P2 band and 41% of the P3 band time. The bad weather we had at the beginning of the period, back in December, really kept us from completing many of the high priority targets in the Sheila field which is now unreachable until the late summer.

There are 10 programs that have been completed out of 27 programs. Some of our users have never been granted HET time before so there is a bit of a learning curve in how to optimize the HET time. Hopefully our Synoptic Supernova teams can make excellent use of the remaining 2 months of the period.

For the next 2 weeks (during the bright time) we will be concentrating on making further improvements to our telescope control systems, wavefront sensor systems and instrument control systems. It is going to be a very busy 2 weeks!

Update from January 9 2017

We have finished another dark period from the first trimester back observing with the HET. Recall that we are only observing during the two weeks around the new moon. Thus, we finished up last week with our targets that can be observed with some moon. We have completed 63.4 hours of charged time with typical setup times of 10 minutes for LRS2 targets. Some programs, particularly those from our smaller partners, have been completed finished while a few of the larger programs have only had a small dent made in their total Telescope Allocation Committee (TAC) allocations.

There are a few fields that seem to be very popular. The Shelia and Cosmos fields are among the most heavily over-subscribed. We will not be able to complete the requested visits to the Shelia fields but if the weather permits and the engineering efforts during the 1st quarter are not to heavy we might try to sneak in a few more targets.

We are developing a hole at the start of the night after the Shelia field sets and will likely have to deal with that in the next few weeks unless some new targets arrive.